the daily cost to rent a truck is $19.95, with an added cost of $.50 per mile. How much would it cost to rent a truck for one day and drive it for 73 miles?

Found 2 solutions by math_tutor2020, mananth:
Answer by math_tutor2020(3817)


Hint:
Evaluate y = 0.50x + 19.95 when x = 73

Answer by mananth(16946)


the daily cost to rent a truck is $19.95, with an added cost of $.50 per mile. How much would it cost to rent a truck for one day and drive it for 73 miles?
Total Cost=Daily Rental Cost+(Miles Driven×Cost per Mile)
Let total cost including mileage be $y
Let number of miles driven =x

daily cost to rent a truck is $19.95
$.50 per mile
y= 19.95+0.5x

y = 19.95+0.5*73= 56.45
To rent the truck for one day and drive it for 73 miles it will cost $56.45